OpenStudy (anonymous):

Help please hurry!!! Graph the ordered pairs (4,-2) and (1, -1) and draw a line through the two points. Which quadrant is untouched by the line? Explain how you found your answer.

OpenStudy (anonymous):

Quadrant two is untouched.

OpenStudy (anonymous):

\[m = \frac{ y _{2} - y _{1} }{ x _{2} - x _{1} }\] \[y - y _{1} = m(x - x _{1})\] use these two formulae to find out how to graph it.